OpenStudy (anonymous):

Find the coordinates of the midpoint of the segment whose endpoints are given. W (-3, -7) and X (-8, -4)

OpenStudy (anonymous):

take the average in each coordinate, i.e. add and divide by 2

OpenStudy (anonymous):

\[\frac{-3-8}{2}=\frac{-11}{2}\] \[\frac{-7-4}{2}=\frac{-11}{2}\]

OpenStudy (anonymous):

your answer is \[\left(-\frac{11}{2},-\frac{11}{2}\right)\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw